OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5

OpenStudy (anonymous):

if you can help send me a message

imqwerty (imqwerty):

yea i can help can u post a screenshot the equations are all messed up

OpenStudy (anonymous):

no there right you have to find the (x,y)

imqwerty (imqwerty):

oh ok i got it now so 1st we gotta find out the equations of both the lines if u have have 2 points (x1,y1) and (x2,y2) then the equation of line is given like this-\[\frac{ y-y_{1} }{ y_{2}-y_{1} }=\frac{ x-x_{1} }{ x_{2}-x_{1} }\] can u use this information to find the equation of the lines?

OpenStudy (anonymous):

did you get my message

OpenStudy (anonymous):

?